OpenStudy (anonymous):

Anita has a collection of sports trading cards. Her collection consists of three basketball cards, six baseball cards, seven hockey cards, and eight football cards. Anita picked one card from her collection.

OpenStudy (anonymous):

whats the question

OpenStudy (anonymous):

What is the probability that Anita picked a football card?

jimthompson5910 (jim_thompson5910):

There are 3+6+7+8 = 24 cards total Of these cards, there are 8 football cards So the probability is... P(Picking Football Card) = (# of football cards)/(# of cards total) P(Picking Football Card) = (8)/(24) P(Picking Football Card) = 1/3

OpenStudy (anonymous):

thankyou (:

jimthompson5910 (jim_thompson5910):

you're welcome